# Chain-of-Thought Unfaithfulness as Disguised Accuracy

### v1

@@ A.1 Implementation Details
As shown in table 2, we don’t evaluate on the entire test set due to resource constraints. Instead, we either use the full dataset or sample 500 examples, whichever is less.
@@
Due to resource constraints, we aren’t able to fit Llama 2 70b or FLAN-UL2 on a single NVIDIA A100 GPU. In order to still get results, we use quantized 4-bit and 8-bit alternatives, respectively. For evaluation settings, these quantized models have shown similar performance as their unquantized counterparts (Dettmers et al., 2024).
Shuffling Answer Choice Order
In Figure 2 we describe the same and different conditions for ordering the answer choices of a MCQ. Another option is to simply use the ordering of the answer choices from the original dataset. We report results under this original condition in Figures 6-7 and Tables 3-4. In practice the original and same conditions yield similar results.
@@
ARC-Challenge FLAN-T5 77m 269 25.85 26.28 64.42
FLAN-T5 248m 269 38.74 38.57 62.2
FLAN-T5 783m 269 42.32 46.76 51.19
FLAN-T5 2.85b 269 64.85 60.15 60.58
FLAN-T5 11.3b 269 25.85 26.28 64.42
FLAN-UL2 20b 269 79.8 66.8 71.6
Llama 2 7b 269 57.59 58.36 66.64
Llama 2 13b 269 63.41 67.0 70.88
Llama 2 70b 269 76.72 76.62 76.42
Pythia DPO 70m 269 23.6 24.0 99.6
Pythia DPO 160m 269 22.7 23.98 82.34
Pythia DPO 410m 269 23.21 22.87 89.33
Pythia DPO 1b 269 21.8 28.0 35.6
Pythia DPO 1.4b 269 26.28 27.13 16.47
Pythia DPO 2.8b 269 23.8 25.8 33.6
ARC-Easy FLAN-T5 77m 500 28.41 25.84 62.04
FLAN-T5 248m 500 53.2 51.18 65.57
FLAN-T5 783m 500 53.49 57.53 54.17
FLAN-T5 2.85b 500 28.41 25.84 62.04
FLAN-T5 11.3b 500 28.41 25.84 62.04
FLAN-UL2 20b 500 93.6 84.2 85.4
Llama 2 7b 500 75.42 72.77 75.93
Llama 2 13b 500 83.52 83.12 82.88
Llama 2 70b 500 92.35 91.8 91.89
Pythia DPO 70m 500 22.8 22.8 100.0
Pythia DPO 160m 500 25.17 25.34 82.53
Pythia DPO 410m 500 25.12 25.45 90.86
Pythia DPO 1b 500 23.2 26.6 39.6
Pythia DPO 1.4b 500 27.6 28.6 31.8
Pythia DPO 2.8b 500 24.4 24.4 34.2
@@
HellaSwag FLAN-T5 77m 268 42.0 31.4 42.6
FLAN-T5 248m 268 42.7 38.41 59.04
FLAN-T5 783m 268 78.26 64.66 70.58
FLAN-T5 2.85b 268 91.04 86.21 89.73
FLAN-T5 11.3b 268 90.54 85.18 87.79
FLAN-UL2 20b 268 83.6 75.2 81.8
Llama 2 7b 268 51.71 44.72 55.28
Llama 2 13b 268 59.49 53.43 60.29
Llama 2 70b 268 72.01 62.69 69.03
Pythia DPO 70m 268 25.13 25.03 98.77
Pythia DPO 160m 268 24.97 25.0 99.52
Pythia DPO 410m 268 25.22 25.22 96.03
Pythia DPO 1b 268 25.38 24.79 34.6
Pythia DPO 1.4b 268 24.82 23.61 32.07
Pythia DPO 2.8b 268 25.52 27.08 24.54
LogiQA FLAN-T5 77m 500 28.26 26.42 83.26
FLAN-T5 248m 500 30.88 31.64 58.22
FLAN-T5 783m 500 28.88 27.19 54.84
FLAN-T5 2.85b 500 31.8 33.49 50.84
FLAN-T5 11.3b 500 37.17 31.95 41.01
FLAN-UL2 20b 500 33.8 31.2 48.4
Llama 2 7b 500 36.41 31.49 51.77
Llama 2 13b 500 39.78 34.41 51.77
Llama 2 70b 500 43.93 41.63 57.91
Pythia DPO 70m 500 20.4 20.2 99.4
Pythia DPO 160m 500 20.12 22.89 77.73
Pythia DPO 410m 500 20.74 21.81 90.48
Pythia DPO 1b 500 26.42 23.81 8.91
Pythia DPO 1.4b 500 24.8 22.6 25.0
Pythia DPO 2.8b 500 29.8 27.8 33.2
MMLU FLAN-T5 77m 252 27.93 26.05 38.3
FLAN-T5 248m 252 33.54 32.53 56.4
FLAN-T5 783m 252 42.19 35.57 44.19
FLAN-T5 2.85b 252 42.2 43.4 46.6
FLAN-T5 11.3b 252 48.0 37.2 40.4
FLAN-UL2 20b 252 50.4 46.4 50.2
Llama 2 7b 252 50.67 52.45 59.04
Llama 2 13b 252 53.81 55.62 63.07
Llama 2 70b 252 54.8 64.4 56.6
Pythia DPO 70m 252 20.8 21.0 99.0
Pythia DPO 160m 252 22.77 22.86 99.5
Pythia DPO 410m 252 22.7 23.0 89.0
Pythia DPO 1b 252 24.21 21.43 35.32
Pythia DPO 1.4b 252 27.85 27.03 33.38
Pythia DPO 2.8b 252 27.27 25.83 33.66
@@
OpenBookQA FLAN-T5 77m 287 27.8 21.6 59.6
FLAN-T5 248m 287 46.0 41.4 67.2
FLAN-T5 783m 287 44.8 47.6 47.2
FLAN-T5 2.85b 287 70.4 65.2 64.0
FLAN-T5 11.3b 287 74.4 33.4 29.2
FLAN-UL2 20b 287 79.6 69.6 71.2
Llama 2 7b 287 59.4 57.2 62.0
Llama 2 13b 287 63.8 65.0 67.2
Llama 2 70b 287 75.2 74.2 73.8
Pythia DPO 70m 287 27.6 28.2 89.6
Pythia DPO 160m 287 27.6 25.4 80.8
Pythia DPO 410m 287 27.6 25.8 91.8
Pythia DPO 1b 287 25.2 27.8 15.0
Pythia DPO 1.4b 287 26.8 23.8 34.8
Pythia DPO 2.8b 287 22.87 22.59 23.69
TruthfulQA FLAN-T5 77m 500 5.88 15.42 70.13
FLAN-T5 248m 500 27.66 12.12 59.0
FLAN-T5 783m 500 8.08 12.36 49.82
FLAN-T5 2.85b 500 8.32 20.93 44.55
FLAN-T5 11.3b 500 23.62 29.62 43.33
FLAN-UL2 20b 500 47.6 35.4 51.6
Llama 2 7b 500 28.89 35.37 58.02
Llama 2 13b 500 24.11 46.39 45.41
Llama 2 70b 500 41.98 45.78 57.04
Pythia DPO 70m 500 100.0 98.2 98.2
Pythia DPO 160m 500 99.88 66.95 66.83
Pythia DPO 410m 500 97.55 82.5 80.66
Pythia DPO 1b 500 10.16 60.59 13.71
Pythia DPO 1.4b 500 0.8 15.0 39.0
Pythia DPO 2.8b 500 10.8 22.6 31.2
@@
Llama 2 70b 254 27.56 37.8 36.61
Pythia DPO 70m 254 19.69 20.87 92.52
Pythia DPO 160m 254 22.83 20.87 59.84
Pythia DPO 410m 254 18.9 18.5 85.43
Pythia DPO 1b 254 15.35 20.08 8.66
Pythia DPO 1.4b 254 20.87 18.5 20.47
Pythia DPO 2.8b 254 23.62 18.9 24.8
ARC-Challenge FLAN-T5 77m 269 24.23 22.27 65.1
FLAN-T5 248m 269 37.12 37.46 62.2
FLAN-T5 783m 269 43.0 46.42 49.91
FLAN-T5 2.85b 269 63.91 61.69 62.29
FLAN-T5 11.3b 269 25.43 25.6 62.2
FLAN-UL2 20b 269 78.8 65.8 69.4
Llama 2 7b 269 56.57 56.14 65.1
Llama 2 13b 269 62.37 63.14 68.6
Llama 2 70b 269 78.29 74.68 76.83
Pythia DPO 70m 269 22.2 22.2 99.8
Pythia DPO 160m 269 25.0 25.2 99.4
Pythia DPO 410m 269 29.37 25.28 85.87
Pythia DPO 1b 269 25.68 25.77 12.71
Pythia DPO 1.4b 269 27.56 24.66 15.53
Pythia DPO 2.8b 269 27.6 25.4 32.2
ARC-Easy FLAN-T5 77m 500 26.89 25.97 62.75
FLAN-T5 248m 500 53.83 49.16 65.24
FLAN-T5 783m 500 53.2 57.7 55.68
FLAN-T5 2.85b 500 27.48 26.85 61.24
FLAN-T5 11.3b 500 28.62 27.9 62.16
FLAN-UL2 20b 500 93.2 84.8 85.0
Llama 2 7b 500 75.97 72.35 75.51
Llama 2 13b 500 84.55 83.86 83.66
Llama 2 70b 500 93.07 90.91 91.09
Pythia DPO 70m 500 24.6 24.4 99.8
Pythia DPO 160m 500 26.3 24.79 81.1
Pythia DPO 410m 500 24.56 24.7 90.99
Pythia DPO 1b 500 28.6 26.8 35.2
Pythia DPO 1.4b 500 26.56 25.0 17.0
Pythia DPO 2.8b 500 28.6 26.4 31.4
@@
HellaSwag FLAN-T5 77m 268 44.4 33.8 40.8
FLAN-T5 248m 268 42.52 38.24 59.91
FLAN-T5 783m 268 78.36 65.02 70.92
FLAN-T5 2.85b 268 90.83 86.32 89.65
FLAN-T5 11.3b 268 90.75 85.11 87.93
FLAN-UL2 20b 268 83.2 72.6 79.2
Llama 2 7b 268 50.34 43.16 55.48
Llama 2 13b 268 57.68 47.78 54.02
Llama 2 70b 268 73.8 57.4 66.0
Pythia DPO 70m 268 24.77 24.84 98.73
Pythia DPO 160m 268 25.33 25.46 99.7
Pythia DPO 410m 268 24.18 24.13 95.51
Pythia DPO 1b 268 24.92 24.12 34.25
Pythia DPO 1.4b 268 23.04 25.05 32.18
Pythia DPO 2.8b 268 26.2 24.4 26.4
LogiQA FLAN-T5 77m 500 24.58 24.58 84.95
FLAN-T5 248m 500 29.49 27.5 56.53
FLAN-T5 783m 500 24.27 27.65 56.22
FLAN-T5 2.85b 500 32.72 32.72 48.39
FLAN-T5 11.3b 500 36.56 31.03 39.78
FLAN-UL2 20b 500 37.0 33.8 48.4
Llama 2 7b 500 34.41 33.64 51.31
Llama 2 13b 500 36.87 33.33 51.31
Llama 2 70b 500 42.09 39.63 58.53
Pythia DPO 70m 500 26.2 26.2 99.8
Pythia DPO 160m 500 24.12 24.73 79.88
Pythia DPO 410m 500 24.58 25.81 88.17
Pythia DPO 1b 500 28.73 21.97 7.37
Pythia DPO 1.4b 500 24.0 23.4 26.4
Pythia DPO 2.8b 500 25.0 23.4 33.0
MMLU FLAN-T5 77m 252 27.87 26.24 38.85
FLAN-T5 248m 252 33.96 33.26 54.67
FLAN-T5 783m 252 39.97 35.31 45.37
FLAN-T5 2.85b 252 39.8 40.0 44.2
FLAN-T5 11.3b 252 49.8 34.0 41.4
FLAN-UL2 20b 252 47.6 44.8 56.8
Llama 2 7b 252 50.17 49.89 59.36
Llama 2 13b 252 54.81 55.59 63.01
Llama 2 70b 252 52.8 64.4 56.2
Pythia DPO 70m 252 29.0 29.0 98.8
Pythia DPO 160m 252 26.46 26.43 99.46
Pythia DPO 410m 252 23.4 25.0 84.4
Pythia DPO 1b 252 25.31 25.43 31.56
Pythia DPO 1.4b 252 25.03 24.21 33.84
Pythia DPO 2.8b 252 26.66 26.33 34.11
@@
OpenBookQA FLAN-T5 77m 287 24.4 25.4 61.4
FLAN-T5 248m 287 28.6 24.4 61.6
FLAN-T5 783m 287 27.2 28.2 61.8
FLAN-T5 2.85b 287 29.6 27.0 58.8
FLAN-T5 11.3b 287 26.2 24.0 59.8
FLAN-UL2 20b 287 79.2 70.4 72.8
Llama 2 7b 287 58.4 55.8 61.2
Llama 2 13b 287 65.0 63.8 65.6
Llama 2 70b 287 80.0 73.4 74.2
Pythia DPO 70m 287 22.0 24.2 91.0
Pythia DPO 160m 287 24.2 24.2 99.8
Pythia DPO 410m 287 25.6 26.6 90.8
Pythia DPO 1b 287 29.4 24.8 14.4
Pythia DPO 1.4b 287 25.8 26.4 32.8
Pythia DPO 2.8b 287 24.74 26.48 26.83
TruthfulQA FLAN-T5 77m 500 19.34 20.81 71.73
FLAN-T5 248m 500 18.97 21.79 71.24
FLAN-T5 783m 500 23.26 23.87 70.62
FLAN-T5 2.85b 500 18.85 19.83 70.75
FLAN-T5 11.3b 500 20.93 20.93 71.6
FLAN-UL2 20b 500 46.2 36.6 55.8
Llama 2 7b 500 28.15 43.21 53.98
Llama 2 13b 500 31.09 47.86 50.92
Llama 2 70b 500 46.35 56.0 60.44
Pythia DPO 70m 500 26.0 26.6 99.2
Pythia DPO 160m 500 23.5 23.87 64.87
Pythia DPO 410m 500 24.0 22.8 82.2
Pythia DPO 1b 500 21.05 22.52 17.87
Pythia DPO 1.4b 500 24.4 23.4 45.8
Pythia DPO 2.8b 500 23.4 24.8 34.4
@@
Pythia DPO 70m 254 22.05 14.96 24.8
Pythia DPO 160m 254 15.35 17.72 24.02
Pythia DPO 410m 254 19.69 20.47 22.05
Pythia DPO 1b 254 22.05 21.26 23.23
Pythia DPO 1.4b 254 24.8 21.65 19.29
@@
ARC-Challenge FLAN-T5 77m 269 24.15 25.6 22.95
FLAN-T5 248m 269 38.82 36.6 52.3
FLAN-T5 783m 269 43.43 42.75 40.7
FLAN-T5 2.85b 269 64.33 59.73 59.73
FLAN-T5 11.3b 269 26.11 26.19 26.19
FLAN-UL2 20b 269 78.6 66.0 67.6
Llama 2 7b 269 55.2 54.44 52.13
Llama 2 13b 269 64.16 65.44 62.54
Llama 2 70b 269 77.13 77.02 75.2
Pythia DPO 70m 269 24.0 26.2 22.8
Pythia DPO 160m 269 25.4 25.0 22.8
Pythia DPO 410m 269 24.2 25.4 23.8
Pythia DPO 1b 269 25.2 25.4 22.8
Pythia DPO 1.4b 269 25.34 23.9 25.05
Pythia DPO 2.8b 269 24.6 28.0 22.0
ARC-Easy FLAN-T5 77m 500 27.06 25.59 25.25
FLAN-T5 248m 500 53.07 49.16 58.96
FLAN-T5 783m 500 53.41 57.45 46.59
FLAN-T5 2.85b 500 24.49 25.34 25.84
FLAN-T5 11.3b 500 26.77 27.61 24.49
FLAN-UL2 20b 500 93.6 83.4 83.6
Llama 2 7b 500 74.71 71.04 67.26
Llama 2 13b 500 81.57 79.04 75.38
Llama 2 70b 500 91.61 91.07 89.18
Pythia DPO 70m 500 24.4 25.8 24.4
Pythia DPO 160m 500 25.13 25.29 25.13
Pythia DPO 410m 500 25.89 24.84 24.89
Pythia DPO 1b 500 22.0 27.2 24.6
Pythia DPO 1.4b 500 24.54 24.07 26.05
Pythia DPO 2.8b 500 26.8 24.4 24.2

### v3

@@ A.1 Implementation Details
As shown in Table 2, we do not always evaluate the entire test sets due to resource constraints. Instead, we either use the full dataset or sample 500 examples, whichever is less.
@@
Due to resource constraints, we are not able to fit Llama 2 70b or FLAN-UL2 on a single NVIDIA A100 GPU. In order to still get results, we use quantized 4-bit and 8-bit alternatives, respectively. For evaluation settings, these quantized models have shown similar performance as their unquantized counterparts (Dettmers et al., 2024).
A.2 Experiments with Different Answer Ordering
In Equation (3) we introduce normalized unfaithfulness which accounts for the known inductive bias in some models to consistently select the same answer choice. Another approach to mitigate this bias is by applying different shuffling strategies and comparing their effect on the unfaithfulness metric. The two strategies we use are described below, and are demonstrated in Figure 5.
Figure 5: Illustration of the same vs. different ordering conditions for MCQs.
One problem with the different condition is that it introduces two treatments at once: the ordering of the answer and the provision of CoT. This means that when we compare results of the unfaithfulness metric from this condition with those from the same condition, we cannot precisely assign which treatment caused the effects. For example, it could be that changes in unfaithfulness in the different condition are due to sensitivity to answer ordering and not CoT. As seen in Fig. 6 the effect of different ordering is empirically similar to that of normalized unfaithfulness. This is also the case in Fig. 7 where the observed correlation with accuracy is even stronger ($R^{2}=0.862$) than for normalized unfaithfulness. Due to its cleaner interpretation, we opt to present our normalized unfaithfulness, but include the results for the different condition here.
Figure 6: Lanham et al. (2023)’s unfaithfulness (left), normalized unfaithfulness (center), and Lanham et al. (2023)’s unfaithfulness when choice ordering is different with- and without- CoT (right). Normalized unfaithfulness shows an V-shape, but normalized unfaithfulness and different-ordering unfaithfulness show a scaling relationship.
Figure 7: Correlation plot showing accuracy vs. unfaithfulness when the prompt choices are shuffled between the CoT and no-CoT conditions. The two metrics are highly correlated ($R^{2}=0.862$).
Figure 8: Given a different ordering of MCQ options, how often each model responds with the same letter (left) and same answer (right). As model size increases, models favor the same answer more and the same letter less.
@@ A.2 Experiments with Different Answer Ordering
ARC-Challenge FLAN-T5 77m 1172 25.85 26.28 64.42
FLAN-T5 248m 1172 38.74 38.57 62.20
FLAN-T5 783m 1172 42.32 46.76 51.19
FLAN-T5 2.85b 1172 64.85 60.15 60.58
FLAN-T5 11.3b 1172 25.85 26.28 64.42
FLAN-T5 20b 500 79.80 66.80 71.60
Llama 2 7b 1172 57.59 58.36 66.64
Llama 2 13b 697 63.41 67.00 70.88
Llama 2 70b 988 76.72 76.62 76.42
Pythia DPO 70m 500 23.60 24.00 99.60
Pythia DPO 160m 1172 22.70 23.98 82.34
Pythia DPO 410m 1172 23.21 22.87 89.33
Pythia DPO 1b 500 21.80 28.00 35.60
Pythia DPO 1.4b 1172 26.28 27.13 16.47
Pythia DPO 2.8b 500 23.80 25.80 33.60
ARC-Easy FLAN-T5 77m 2376 28.41 25.84 62.04
FLAN-T5 248m 2376 53.20 51.18 65.57
FLAN-T5 783m 2376 53.49 57.53 54.17
FLAN-T5 2.85b 2376 28.41 25.84 62.04
FLAN-T5 11.3b 2376 28.41 25.84 62.04
FLAN-T5 20b 500 93.60 84.20 85.40
Llama 2 7b 2376 75.42 72.77 75.93
Llama 2 13b 1250 83.52 83.12 82.88
Llama 2 70b 1098 92.35 91.80 91.89
Pythia DPO 70m 500 22.80 22.80 100.00
Pythia DPO 160m 2376 25.17 25.34 82.53
Pythia DPO 410m 2122 25.12 25.45 90.86
Pythia DPO 1b 500 23.20 26.60 39.60
Pythia DPO 1.4b 500 27.60 28.60 31.80
Pythia DPO 2.8b 500 24.40 24.40 34.20
@@
HellaSwag FLAN-T5 77m 500 42.00 31.40 42.60
FLAN-T5 248m 10042 42.70 38.41 59.04
FLAN-T5 783m 10042 78.26 64.66 70.58
FLAN-T5 2.85b 10042 91.04 86.21 89.73
FLAN-T5 11.3b 10042 90.54 85.18 87.79
FLAN-T5 20b 500 83.60 75.20 81.80
Llama 2 7b 4434 51.71 44.72 55.28
Llama 2 13b 627 59.49 53.43 60.29
Llama 2 70b 500 72.01 62.69 69.03
Pythia DPO 70m 6891 25.13 25.03 98.77
Pythia DPO 160m 3940 24.97 25.00 99.52
Pythia DPO 410m 2117 25.22 25.22 96.03
Pythia DPO 1b 2731 25.38 24.79 34.60
Pythia DPO 1.4b 2236 24.82 23.61 32.07
Pythia DPO 2.8b 1536 25.52 27.08 24.54
LogiQA FLAN-T5 77m 651 28.26 26.42 83.26
FLAN-T5 248m 651 30.88 31.64 58.22
FLAN-T5 783m 651 28.88 27.19 54.84
FLAN-T5 2.85b 651 31.80 33.49 50.84
FLAN-T5 11.3b 651 37.17 31.95 41.01
FLAN-T5 20b 500 33.80 31.20 48.40
Llama 2 7b 651 36.41 31.49 51.77
Llama 2 13b 651 39.78 34.41 51.77
Llama 2 70b 651 43.93 41.63 57.91
Pythia DPO 70m 500 20.40 20.20 99.40
Pythia DPO 160m 651 20.12 22.89 77.73
Pythia DPO 410m 651 20.74 21.81 90.48
Pythia DPO 1b 651 26.42 23.81 8.91
Pythia DPO 1.4b 500 24.80 22.60 25.00
Pythia DPO 2.8b 500 29.80 27.80 33.20
MMLU FLAN-T5 77m 13985 27.93 26.05 38.30
FLAN-T5 248m 13985 33.54 32.53 56.40
FLAN-T5 783m 6379 42.19 35.57 44.19
FLAN-T5 2.85b 500 42.20 43.40 46.60
FLAN-T5 11.3b 500 48.00 37.20 40.40
FLAN-T5 20b 500 50.40 46.40 50.20
Llama 2 7b 4766 50.67 52.45 59.04
Llama 2 13b 3520 53.81 55.62 63.07
Llama 2 70b 500 54.80 64.40 56.60
Pythia DPO 70m 500 20.80 21.00 99.00
Pythia DPO 160m 2213 22.77 22.86 99.50
Pythia DPO 410m 1991 22.70 23.00 89.00
Pythia DPO 1b 500 24.21 21.43 35.32
Pythia DPO 1.4b 2172 27.85 27.03 33.38
Pythia DPO 2.8b 1533 27.27 25.83 33.66
@@
OpenBookQA FLAN-T5 77m 500 27.80 21.60 59.60
FLAN-T5 248m 500 46.00 41.40 67.20
FLAN-T5 783m 500 44.80 47.60 47.20
FLAN-T5 2.85b 500 70.40 65.20 64.00
FLAN-T5 11.3b 500 74.40 33.40 29.20
FLAN-T5 20b 500 79.60 69.60 71.20
Llama 2 7b 500 59.40 57.20 62.00
Llama 2 13b 500 63.80 65.00 67.20
Llama 2 70b 500 75.20 74.20 73.80
Pythia DPO 70m 500 27.60 28.20 89.60
Pythia DPO 160m 500 27.60 25.40 80.80
Pythia DPO 410m 500 27.60 25.80 91.80
Pythia DPO 1b 500 25.20 27.80 15.00
Pythia DPO 1.4b 500 26.80 23.80 34.80
Pythia DPO 2.8b 500 22.87 22.59 23.69
TruthfulQA FLAN-T5 77m 817 5.88 15.42 70.13
FLAN-T5 248m 817 27.66 12.12 59.00
FLAN-T5 783m 817 8.08 12.36 49.82
FLAN-T5 2.85b 817 8.32 20.93 44.55
FLAN-T5 11.3b 817 23.62 29.62 43.33
FLAN-T5 20b 500 47.60 35.40 51.60
Llama 2 7b 817 28.89 35.37 58.02
Llama 2 13b 817 24.11 46.39 45.41
Llama 2 70b 817 41.98 45.78 57.04
Pythia DPO 70m 500 100.00 98.20 98.20
Pythia DPO 160m 817 99.88 66.95 66.83
Pythia DPO 410m 817 97.55 82.50 80.66
Pythia DPO 1b 817 10.16 60.59 13.71
Pythia DPO 1.4b 500 0.80 15.00 39.00
Pythia DPO 2.8b 500 10.80 22.60 31.20
@@
Llama 2 70b 254 27.56 37.80 36.61
Pythia DPO 70m 254 19.69 20.08 99.61
Pythia DPO 160m 254 24.41 23.62 99.21
Pythia DPO 410m 254 18.91 18.57 85.43
Pythia DPO 1b 254 25.59 19.69 42.13
Pythia DPO 1.4b 254 20.87 18.54 20.47
Pythia DPO 2.8b 254 20.47 22.83 20.08
ARC-Challenge FLAN-T5 77m 1172 24.23 22.27 65.10
FLAN-T5 248m 1172 37.12 37.46 62.20
FLAN-T5 783m 1172 43.00 46.42 49.91
FLAN-T5 2.85b 1172 63.91 61.69 62.29
FLAN-T5 11.3b 1172 25.43 25.60 62.20
FLAN-T5 20b 500 78.80 65.80 69.40
Llama 2 7b 1172 56.57 56.14 65.10
Llama 2 13b 1172 62.37 63.14 68.60
Llama 2 70b 1027 78.29 74.68 76.83
Pythia DPO 70m 500 24.60 24.80 99.80
Pythia DPO 160m 500 24.85 25.44 99.41
Pythia DPO 410m 500 25.00 25.40 84.00
Pythia DPO 1b 500 29.20 25.40 38.20
Pythia DPO 1.4b 389 24.42 22.11 32.39
Pythia DPO 2.8b 500 27.60 25.40 32.20
ARC-Easy FLAN-T5 77m 2376 26.89 25.97 62.75
FLAN-T5 248m 2376 53.83 49.16 65.24
FLAN-T5 783m 2376 53.20 57.70 55.68
FLAN-T5 2.85b 2376 27.48 26.85 61.24
FLAN-T5 11.3b 2376 28.62 27.90 62.16
FLAN-T5 20b 500 93.20 84.80 85.00
Llama 2 7b 2376 75.97 72.35 75.51
Llama 2 13b 1010 84.55 83.86 83.66
Llama 2 70b $1$ 93.07 90.91 91.09
Pythia DPO 70m 500 25.60 25.60 99.43
Pythia DPO 160m 500 28.60 28.60 98.87
Pythia DPO 410m 500 24.20 24.80 83.84
Pythia DPO 1b 500 27.60 23.40 33.60
Pythia DPO 1.4b 500 26.00 24.20 30.64
Pythia DPO 2.8b 500 28.60 26.40 31.43
@@
HellaSwag FLAN-T5 77m 500 44.40 33.80 40.80
FLAN-T5 248m 10042 42.52 38.24 59.91
FLAN-T5 783m 10042 78.36 65.02 70.92
FLAN-T5 2.85b 10042 90.83 86.32 89.65
FLAN-T5 11.3b 10042 90.75 85.11 87.93
FLAN-T5 20b 500 83.20 72.60 79.20
Llama 2 7b 4557 50.34 43.16 55.48
Llama 2 13b 3847 57.68 47.78 54.02
Llama 2 70b 500 73.80 57.40 66.00
Pythia DPO 70m 500 27.60 27.40 98.20
Pythia DPO 160m 500 24.60 24.40 99.60
Pythia DPO 410m 500 24.63 24.63 91.71
Pythia DPO 1b 500 23.00 25.80 31.00
Pythia DPO 1.4b 500 24.20 26.20 37.00
Pythia DPO 2.8b 500 25.40 21.80 22.00
LogiQA FLAN-T5 77m 651 24.58 24.58 84.95
FLAN-T5 248m 651 29.49 27.50 56.53
FLAN-T5 783m 651 24.27 27.65 56.22
FLAN-T5 2.85b 651 32.72 32.72 48.39
FLAN-T5 11.3b 651 36.56 31.03 39.78
FLAN-T5 20b 500 37.00 33.80 48.40
Llama 2 7b 651 34.41 33.64 51.31
Llama 2 13b 651 36.87 33.33 51.31
Llama 2 70b 651 42.09 39.63 58.53
Pythia DPO 70m 500 26.20 26.20 99.80
Pythia DPO 160m 500 25.40 25.40 99.80
Pythia DPO 410m 500 28.60 29.20 78.80
Pythia DPO 1b 500 19.70 25.12 33.00
Pythia DPO 1.4b 500 24.00 23.40 26.40
Pythia DPO 2.8b 500 25.00 23.40 33.00
MMLU FLAN-T5 77m 13985 27.87 26.24 38.85
FLAN-T5 248m 13985 33.96 33.26 54.67
FLAN-T5 783m 6497 39.97 35.31 45.37
FLAN-T5 2.85b 500 39.80 40.00 44.20
FLAN-T5 11.3b 500 49.80 34.00 41.40
FLAN-T5 20b 500 47.60 44.80 56.80
Llama 2 7b 4901 50.17 49.89 59.36
Llama 2 13b 3760 54.81 55.59 63.01
Llama 2 70b 500 52.80 64.40 56.20
Pythia DPO 70m 500 29.00 29.00 98.80
Pythia DPO 160m 3904 26.46 26.43 99.46
Pythia DPO 410m 500 23.40 25.00 84.40
Pythia DPO 1b 2627 25.31 25.43 31.56
Pythia DPO 1.4b 2305 25.03 24.21 33.84
Pythia DPO 2.8b 1504 26.66 26.33 34.11
@@
OpenBookQA FLAN-T5 77m 500 24.40 25.40 61.40
FLAN-T5 248m 500 28.60 24.40 61.60
FLAN-T5 783m 500 27.20 28.20 61.80
FLAN-T5 2.85b 500 29.60 27.00 58.80
FLAN-T5 11.3b 500 26.20 24.00 59.80
FLAN-T5 20b 500 79.20 70.40 72.80
Llama 2 7b 500 58.40 55.80 61.20
Llama 2 13b 500 65.00 63.80 65.60
Llama 2 70b 500 80.00 73.40 74.20
Pythia DPO 70m 500 24.60 25.20 99.40
Pythia DPO 160m 500 24.20 24.20 99.80
Pythia DPO 410m 500 25.60 26.60 90.80
Pythia DPO 1b 500 24.00 22.60 34.80
Pythia DPO 1.4b 500 25.80 26.40 32.80
Pythia DPO 2.8b 500 24.74 26.48 26.83
TruthfulQA FLAN-T5 77m 817 19.34 20.81 71.73
FLAN-T5 248m 817 18.97 21.79 71.24
FLAN-T5 783m 817 23.26 23.87 70.62
FLAN-T5 2.85b 817 18.85 19.83 70.75
FLAN-T5 11.3b 817 20.93 20.93 71.60
FLAN-T5 20b 500 46.20 36.60 55.80
Llama 2 7b 817 28.15 43.21 53.98
Llama 2 13b 817 31.09 47.86 50.92
Llama 2 70b 809 46.35 56.00 60.44
Pythia DPO 70m 500 26.00 26.60 99.20
Pythia DPO 160m 500 21.40 21.60 99.20
Pythia DPO 410m 500 24.00 22.80 82.20
Pythia DPO 1b 500 23.60 21.20 40.80
Pythia DPO 1.4b 500 24.40 23.40 45.80
Pythia DPO 2.8b 500 23.40 24.80 34.40
@@
Pythia DPO 70m 254 18.90 19.29 24.02
Pythia DPO 160m 254 16.54 19.29 16.14
Pythia DPO 410m 254 21.26 25.59 21.26
Pythia DPO 1b 254 21.65 21.65 18.90
Pythia DPO 1.4b 254 20.08 20.08 20.87
@@
ARC-Challenge FLAN-T5 77m 1172 24.15 25.60 22.95
FLAN-T5 248m 1172 38.82 36.60 52.30
FLAN-T5 783m 1172 43.43 42.75 40.70
FLAN-T5 2.85b 1172 64.33 59.73 59.73
FLAN-T5 11.3b 1172 26.11 26.19 26.19
FLAN-T5 20b 500 78.60 66.00 67.60
Llama 2 7b 1172 55.20 54.44 52.13
Llama 2 13b 1172 64.16 65.44 62.54
Llama 2 70b 988 77.13 77.02 75.20
Pythia DPO 70m 500 25.00 26.20 23.00
Pythia DPO 160m 500 25.42 25.06 22.81
Pythia DPO 410m 500 26.00 23.40 24.60
Pythia DPO 1b 500 25.20 25.40 22.80
Pythia DPO 1.4b 500 22.81 26.32 25.00
Pythia DPO 2.8b 500 24.60 28.00 22.00
ARC-Easy FLAN-T5 77m 2376 27.06 25.59 25.25
FLAN-T5 248m 2376 53.07 49.16 58.96
FLAN-T5 783m 2376 53.41 57.45 46.59
FLAN-T5 2.85b 2376 24.49 25.34 25.84
FLAN-T5 11.3b 2376 26.77 27.61 24.49
FLAN-T5 20b 500 93.60 83.40 83.60
Llama 2 7b 2376 74.71 71.04 67.26
Llama 2 13b 2376 81.57 79.04 75.38
Llama 2 70b 1109 91.61 91.07 89.18
Pythia DPO 70m 500 25.40 26.40 25.00
Pythia DPO 160m 500 22.87 26.52 22.63
Pythia DPO 410m 500 24.70 29.17 26.49
Pythia DPO 1b 500 26.00 23.20 23.00
Pythia DPO 1.4b 500 25.40 25.20 27.00
Pythia DPO 2.8b 500 26.80 24.40 24.20